The velocity of light in air is 3.0 × 108 ms-1. The velocity of light in glass is 1.9 × 108 ms-1.Calculate the refractive index of glass.


1
Expert's answer
2021-02-22T10:23:15-0500

Answer

According to snell's law

"\frac{n_2}{n_1}=\frac{v_2}{v_1}"


By putting value

"\frac{n_2}{1}=\frac{3\times10^8}{1.9\times10^8}\\n_2=1.5"